solve algebraically: 6x-3y+8z=4
-x+9y-2z=2
4x+6y-4z=3

A)(5/2, 5/3, -3/4)
B)(1, -2, -1)
C)(1/2, 1/3, 1/4)
D)(5, -2, 5/4)

Answer by stanbon(75887)

solve algebraically:
6x-3y+8z=4
-x+9y-2z=2
4x+6y-4z=3
A)(5/2, 5/3, -3/4)
B)(1, -2, -1)
C)(1/2, 1/3, 1/4)
D)(5, -2, 5/4)
-----------------------------------
I used a matrix solution method to get:
x = 1/2
y = 1/3
z = 1/4